# A Dialogue Dataset for Situated PragMATic Reasoning

### v1

@@ Abstract
Pragmatic reasoning aims at resolving implicit meanings that commonly occur in real-life and is crucial for building communicative social agents. We introduce a new benchmark, Diplomat, aiming at a unified paradigm for pragmatic reasoning and situated conversational understanding. Compared with previous works that treat different figurative expressions (e.g., metaphor, sarcasm) as individual tasks, Diplomat provides a unified understanding towards general pragmatic understanding. Our dataset is created using Amazon Mechanical Turk (AMT), resulting in $4,177$ multi-turn dialogues. In company with the dataset, we propose two tasks: Pragmatic Identification and Reasoning and Conversational Question Answering. Experimental results with state-of-the-art (SOTA) neural architectures demonstrate that: 1) large language models show poor performances in this subjective topic. 2) Context understanding is a crucial factor in building benign human-machine interaction. 3) Current models defect in the application of pragmatic reasoning. As a result, we call on more attention to improve the ability of context understanding, reasoning and implied meaning modeling.
@@ 1 Introduction
In order to step towards a general human-like communicative agent, in this work, we introduce Diplomat, a real-life conversational dataset that focuses on pragmatic reasoning. Diplomat stems from an interview dataset Majumder et al. (2020), and experiences three steps of curation: automatic selection, fine-grained manual annotation and human refinement (Sec. 3). Our dataset owns $4,177$ dialogues and covers a vocabulary of $48,900$ words. More than that, human-annotated answers reach the amount of $6,494$ and hold a vocabulary size of $20,000$. Along with the dataset, we propose two tasks, Pragmatic Identification and Reasoning (PIR) and Conversational Question Answering (CQA), to benchmark machines’ pragmatic reasoning capabilities (Sec. 4). The CQA task possesses $19,482$ questions concerning the content of collected dialogues and the answers to the questions are written by humans. We run extensive experiments on previous state-of-the-art models on Diplomat (Sec. 5). The best model achieves less than 0.70 accuracy score in PIR, and none of the models achieve more than 0.50 accuracy score for CQA. Moreover, we test previous pre-trained LLMs’ (including ChatGPT) zero-shot reasoning capability with a natural language inference (NLI) task. Regarding the experimental results provided, the significance of pragmatic reasoning speaks for itself. Throughout a thorough analysis of the limitations of current models, we aim to shed light on future research toward building general conversational agents.
@@ B.2 Details For Fine-grained Annotation
AMT is integral to our process. To ensure clarity and consistency, we provide explicit instructions to the workers. Additionally, to further elucidate the objectives of our study, we offer illustrative examples. The task itself is presented below the instructions and examples, with the dialogue provided and corresponding turn numbers for workers to select. Furthermore, as workers check a checkbox, we prompt them to select a confidence score and provide a rationale. In order to strike a balance between our budget, the quality of annotations, and the speed of annotation, we have determined the compensation of \$0.1 per completed task. The whole view of the worker interface is presented in  Figure 6. After the annotation process, we collect responses that are assigned with a confidence score of 4 or higher.
@@ B.3 Details on Human Refinements
Disturbing choices are chosen based on the BERTScore metric  Zhang et al. (2020b). The rationale with the highest similarity, as determined by other dialogues, is selected and included in the pool of candidate options. The instructions provided to the workers align with those used for Fine-grained Annotation, wherein they are also instructed to assign a confidence score to their responses. The remuneration for workers is set at \$0.05 per task. Also, the worker interface is included in  Figure 7.
@@
In order to guarantee the quality of annotated data, the qualification rules for workers are strict and can be found in  Table 7
@@ C.1 Pragmatic Identification and Reasoning (PIR)
In this section, we will offer details on the experiment results. The PIR task encompasses three distinct settings: C normal-$\rightarrow$ P, CP normal-$\rightarrow$ R, C normal-$\rightarrow$ PR. In the C normal-$\rightarrow$ P setting, models are trained for 20 epochs, employing a batch size as indicated in  Table 9, a learning rate of $52e-$ 5, and weight decay of $0.01$. As for CP normal-$\rightarrow$ R , the hyperparameters adopted are listed in  Table 8. For the C normal-$\rightarrow$ PR setting, there is no training required; instead, we simply load the best checkpoint obtained from the previous training for this task. The concrete implementation is as follows: we initially flatten the test dataset of C normal-$\rightarrow$ P, ensuring that each instance contains both a dialogue and a pragmatic turn extracted from the same dialogue. As for the test dataset of CP normal-$\rightarrow$ R, no modifications are made. It should be noted that, following the processing steps, both datasets own the same dialogues and corresponding pragmatic turns, resulting in identical instance numbers. For an instance to be deemed correct, the models must successfully accomplish both component tasks i.e. succeed in Identification and Reasoning.
Table 8: Batch Size For Models on C normal-$\rightarrow$ P
@@ C.2 Conversational Question Answering (CQA)
ChatGPT was instructed to generate questions for our tasks. The prompt template that starts the questions with "Which" is depicted in  Table 11. Through this methodology, we collected a total of 19,482 questions. To ensure the reliability of the answers provided to these questions, AMT is utilized. The task template is demonstrated in  Figure 8. In our experiment, the hyperparameters adopted are illustrated in  Table 12. o assess the performance of ChatGPT, we conducted testing using the template outlined in Table 14.
@@
Table 14: ChatGPT Test Template of Zero-Shot Natural Language Inference

### v2

@@ Abstract
Pragmatic reasoning plays a pivotal role in deciphering implicit meanings that frequently arise in real-life conversations and is essential for the development of communicative social agents. In this paper, we introduce a novel challenge, DiPlomat, aiming at benchmarking machines’ capabilities on pragmatic reasoning and situated conversational understanding. Compared with previous works that treat different figurative expressions (e.g. metaphor, sarcasm) as individual tasks, DiPlomat provides a cohesive framework towards general pragmatic understanding. Our dataset is created through the utilization of Amazon Mechanical Turk (AMT), resulting in a total of $4,177$ multi-turn dialogues. In conjunction with the dataset, we propose two tasks, Pragmatic Identification and Reasoning (PIR) and Conversational Question Answering (CQA). Experimental results with state-of-the-art (SOTA) neural architectures reveal several significant findings: 1) large language models exhibit poor performance in tackling this subjective domain; 2) comprehensive comprehension of context emerges as a critical factor for establishing benign human-machine interactions; 3) current models defect in the application of pragmatic reasoning. As a result, we call on more attention to improve the ability of context understanding, reasoning, and implied meaning modeling.
@@ 1 Introduction
In order to step towards a general human-like communicative agent, in this work, we introduce DiPlomat, a real-life Dialogue dataset that focuses on Pragmatic reasoning. DiPlomat stems from an interview dataset Majumder et al. (2020), and experiences three steps of curation: automatic selection, fine-grained manual annotation and human refinement (Sec. 3). Our dataset owns $4,177$ dialogues and covers a vocabulary of $48,900$ words. More than that, human-annotated answers reach the amount of $6,494$ and hold a vocabulary size of $20,000$. Along with the dataset, we propose two tasks, Pragmatic Identification and Reasoning (PIR) and Conversational Question Answering (CQA), to benchmark machines’ pragmatic reasoning capabilities (Sec. 4). The CQA task possesses $19,482$ questions concerning the content of collected dialogues and the answers to the questions are written by humans. We run extensive experiments on previous state-of-the-art models on DiPlomat (Sec. 5). The best model achieves less than 0.70 accuracy score in PIR, and none of the models achieve more than 0.50 accuracy score for CQA. Moreover, we test previous pre-trained LLMs’ (including ChatGPT) zero-shot reasoning capability with a natural language inference (NLI) task. Regarding the experimental results provided, the significance of pragmatic reasoning speaks for itself. Throughout a thorough analysis of the limitations of current models, we aim to shed light on future research toward building general conversational agents.
@@ B.2 Details For Fine-grained Annotation
AMT is integral to our process. To ensure clarity and consistency, we provide explicit instructions to the workers. Additionally, to further elucidate the objectives of our study, we offer illustrative examples. The task itself is presented below the instructions and examples, with the dialogue and corresponding turn numbers provided for workers to select. Furthermore, as workers check a checkbox, we prompt them to select a confidence score and provide a rationale. In order to strike a balance between our budget, the quality of annotations, and the speed of annotation, we have determined the compensation of \$0.1 per completed task. The whole view of the worker interface is presented in  Figure 6. After the annotation process, we collect responses that are assigned with a confidence score of 4 or higher.
@@ B.3 Details on Human Refinements
Disturbing choices are chosen based on the BERTScore metric  Zhang et al. (2020b). The rationale with the highest similarity, as determined by other dialogues, is selected and included in the pool of candidate options. The instructions provided to the workers align with those used for Fine-grained Annotation, wherein they are also instructed to assign a confidence score to their responses. The remuneration for workers is set at \$0.05 per task. The worker interface is included in  Figure 7.
@@
In order to guarantee the quality of annotated data, the qualification rules for workers are strict and can be found in  Table 7.
@@ C.1 Pragmatic Identification and Reasoning (PIR)
The PIR task encompasses three distinct settings: C normal-$\rightarrow$ P, CP normal-$\rightarrow$ R, and C normal-$\rightarrow$ PR. In the C normal-$\rightarrow$ P setting, models are trained for 20 epochs, employing a batch size as indicated in Table 9, a learning rate of $52e-$ 5, and weight decay of $0.01$. As for CP normal-$\rightarrow$ R , the hyperparameters adopted are listed in  Table 8. For the C normal-$\rightarrow$ PR setting, there is no training required; instead, we simply load the best checkpoint obtained from the previous training for this task. The concrete implementation is as follows: we initially flatten the test dataset of C normal-$\rightarrow$ P, ensuring that each instance contains both a dialogue and a pragmatic turn extracted from the same dialogue. As for the test dataset of CP normal-$\rightarrow$ R, no modifications are made. It should be noted that, following the processing steps, both datasets own the same dialogues and corresponding pragmatic turns, resulting in identical instance numbers. For an instance to be deemed correct, the models must successfully accomplish both component tasks i.e. succeed in Identification and Reasoning.
Table 8: Hyperparameters for models on CP normal-$\rightarrow$ R
@@ C.2 Conversational Question Answering (CQA)
ChatGPT was instructed to generate questions for our tasks. The prompt template that starts the questions with "Which" is depicted in  Table 11. Through this methodology, we collected a total of 19,482 questions. To ensure the reliability of the answers provided to these questions, AMT is utilized. The task template is demonstrated in  Figure 8. In our experiment, the hyperparameters adopted are illustrated in  Table 12. To assess the performance of ChatGPT, we conducted testing using the template outlined in Table 14.
@@
Table 13: Test ChatGPT: answer questions with only one word.
@@
Following is the question: {question}
Table 14: ChatGPT test template of Zero-Shot CoT